The consumption rate of a certain item in a company is 100 units/day. The ordering
cost per order is ` 100 and the average daily holding cost per unit is 2 paise.
Determine the economic order quantity for placing an order. Also, find the total cost
per unit time (TCU) and the cycle time
1
Expert's answer
2019-03-12T09:35:23-0400

D = 100 units/day, S = 100, H = 0.02.

The economic order quantity for placing an order is:

"EOQ = (2*D*S\/H)^{0.5}" ,

"EOQ = (2*100*100\/0.02)^{0.5}" = 1000 units.

The total cost per unit time is

TCU = D*S/EOQ + 0.5*H*EOQ = 100*100/1000 + 0.5*0.02*1000 = 20.

The cycle time is t = 365/(D/EOQ) = 3650.